OpenStudy (anonymous):

A certain substance has a heat of vaporization of 65.25 kJ/mol. At what Kelvin temperature will the vapor pressure be 6.00 times higher than it was at 319 K? please help

OpenStudy (wolfe8):

So P/T is constant so you have: \[\frac{ P _{1} }{ T _{1} }=\frac{ P _{2} }{ T _{2} }\] and you have P1, T1, and P2=6P1 and you are looking for T2

OpenStudy (anonymous):

ummm could you write that better...it doesn't show in equation form for me

OpenStudy (wolfe8):

Oh maybe it's not working for you. So I wrote P1/T1=P2/T2

OpenStudy (anonymous):

oh ok so it will be T2=6p1/(T1P2)

OpenStudy (anonymous):

also, i dont have the pressure

OpenStudy (wolfe8):

Substitute P2 with 6P1. You will end up cancelling the P :)

OpenStudy (petiteme):

Go Wolfe! :))

OpenStudy (anonymous):

so it will be t2=1/6t1

OpenStudy (wolfe8):

Hmm something looks weird. Here's what I did: P1/T1=P2/T2 (P1/T1)T2=P2 T2=P2(T1/P1) T2=6P1(T1/P1) T2=6T1

OpenStudy (anonymous):

i got it

OpenStudy (wolfe8):

Alright good job.